OpenStudy (mrnebula):

HELP! Divide (x^2 - 4x + 6) / (x - 2) Im confused it says divide but the answer I get isn't any of the options? @zepdrix @Mehek14 @Loser66 @AlexandervonHumboldt2 @ChantySquirrel1129**

OpenStudy (mrnebula):

On this question, like I said I divided but what I get isn't one of my choices. Do you know what I need to do to solve this?

OpenStudy (mrnebula):

x - 2 x - 2 + \[\frac{ 2 }{ x -2 }\] x - 6 + \[\frac{ 8 }{ x - 2 }\] x - 6 + \[\frac{ 16 }{ x - 2 }\]

OpenStudy (mrnebula):

That got entered weird but these r my options

OpenStudy (mathmale):